# Projecting Molecules into Synthesizable Chemical Spaces

### v1

@@ Abstract
Discovering new drug molecules is a pivotal yet challenging process due to the near-infinitely large chemical space and notorious demands on time and resources. Numerous generative models have recently been introduced to accelerate the drug discovery process, but their progression to experimental validation remains limited, largely due to a lack of consideration for synthetic accessibility in practical settings. In this work, we introduce a novel framework that is capable of generating new chemical structures while ensuring synthetic accessibility. Specifically, we introduce a postfix notation of synthetic pathways to represent molecules in chemical space. Then, we design a transformer-based model to translate molecular graphs into postfix notations of synthesis. We highlight the model’s ability to: (a) perform bottom-up synthesis planning more accurately, (b) generate structurally similar, synthesizable analogs for unsynthesizable molecules proposed by generative models with their properties preserved, and (c) explore the local synthesizable chemical space around hit molecules.
@@ 1 Introduction
Designing novel drug molecules, a process that aims to identify molecular structures with desired properties, is known to be notoriously complex, time- and resource-consuming (Wouters et al., 2020). Methods that could accelerate the discovery process are thus of significant interest in medicinal science and the pharmaceutical industry. Historically, two principal computational methodologies have been employed in drug design: virtual screening (Shoichet, 2004) and de novo design (Hartenfeller & Schneider, 2011). Virtual screening involves predicting properties for a predefined set of molecules, known as virtual libraries, and identifying the most promising candidates for subsequent validation. These libraries are usually commercial catalogs (make-on-demand compounds) from chemical vendors (Lyu et al., 2019) and generally consist of molecules with assured synthesizability (Enamine, 2023). However, screening large numbers of molecules is a time-intensive process. Moreover, even libraries on the scale of billions present only a tiny fraction of the theoretical chemical space (Figure 1a), which is nearly infinite in the number of potential synthesizable small molecules (Bohacek et al., 1996). This means that many potentially valuable molecules might not be included in these libraries. Additionally, as the size of the library increases, so does the cost of screening, which in turn limits the effectiveness and efficiency of virtual screening in the drug discovery process.
@@
De novo molecular design methods, unlike virtual screening, dynamically create molecules tailored to specific objectives like protein binding and bioactivity. Early methods of this kind typically adopt combinatorial optimization (Venkatasubramanian et al., 1995; Jensen, 2019; Brown et al., 2019), while recently, neural network enhanced optimization (You et al., 2018; Fu et al., 2021) and generative modeling (Gómez-Bombarelli et al., 2018; Jin et al., 2018) have become popular. These methods aren’t constrained by the limitations of virtual libraries, enabling the efficient exploration of novel chemical entities (see Figure 1c). However, a notable defect with most existing generative models is their tendency to propose synthetically infeasible molecules (Figure 1c) (Gao & Coley, 2020), a concern particularly pronounced in structure-based drug design, an area of increasing interest (Bohacek et al., 1996; Luo et al., 2021; Fu et al., 2022). This issue stems from the inherent ignorance of synthetic accessibility in formulations of most existing generative algorithms, in which molecules are constructed using SMILES strings or atoms and motifs as fundamental units (Gao et al., 2022). This often leads to chemical structures that fall outside synthesizable chemical spaces, posing practical challenges in the experimental validation as well as production of such designs.
@@
Incorporating synthesizability as an additional optimization criterion is a common strategy, yet it introduces complex challenges. Synthesizability is difficult to define precisely due to its sensitivity to subtle structural differences, reaction selectivity, and building block availability inherent in chemical reactions (Gao & Coley, 2020). While notable efforts have been made to estimate synthetic accessibility (Ertl & Schuffenhauer, 2009; Thakkar et al., 2021; Voršilák et al., 2020), reliable quantification of synthetic accessibility is still a far goal. A more effective strategy involves constraining the design process to only consider molecules that are synthesizable, ensuring that algorithms focus exclusively on viable structures. Based on this idea, several models have been introduced (Bradshaw et al., 2019, 2020; Gao et al., 2021; Swanson et al., 2023), though these models have notable limitations, such as poor coverage of synthesizable chemical space or inefficiency in optimizing the properties, as further discussed in Section 2.
@@
To address the recognized challenges, we introduce a framework that learns to “project” molecules into synthesizable chemical spaces (Figure 1d). Given molecules without synthesizability guarantee, our model is able to “fix” them by identifying synthesizable analogs that are structurally similar, thus preserving the structural factors that contributed to the desired properties (Janela & Bajorath, 2022). We represent molecules by a scalable linear representation for synthetic pathways based on postfix notations. By generating synthetic pathways in the form of postfix notations rather than molecular graphs or strings, we can ensure the designed molecules are derivable from purchasable chemical building blocks and expert-defined chemical reaction rules, thereby guaranteeing the synthesizability of resulting molecules. We adopt the transformer architecture (Vaswani et al., 2017), with an encoder for molecular graphs and a decoder to generate postfix notations of synthesis.
@@ 2 Related Work
Combinatorial optimization algorithms use a set of chemical building blocks and reaction templates to make the chemical space combinatorial, and then apply combinatorial optimization algorithms, such as genetic algorithms (Gao et al., 2021), Monte Carlo tree search (Swanson et al., 2023), and Bayesian optimization (Korovina et al., 2020), to search for desired molecules (Figure 1b). Early representative work before the advent of molecular deep learning includes SYNOPSIS (Vinkers et al., 2003) and DOGS (Hartenfeller et al., 2012), which perform virtual reactions to generate candidate molecules and select candidates according to scoring functions. In recent years, advances in deep learning for chemical reactions (Coley et al., 2019a; Schwaller et al., 2019) have inspired methods such as MoleculeChef (Bradshaw et al., 2019), ChemBO (Korovina et al., 2020), and DoGs (Bradshaw et al., 2020) that use neural networks to predict reaction outcomes in a “template-free” manner.
@@
Reinforcement learning approaches, exemplified by PGFS (Gottipati et al., 2020) and REACTOR (Horwood & Noutahi, 2020), formulate the forward synthesis process as Markov decision processes and train agents to explore the chemical space under the guidance of reward functions that quantify desired properties. The agent is provided with a set of reactants and reaction templates and learns to synthesize desirable molecules with them. As the agent is aware of the objective, its search process is directed and expected to be more efficient.
@@
Autoencoder-based methods, represented by MoleculeChef (Bradshaw et al., 2019) and DoGs (Bradshaw et al., 2020), leverage the autoencoder architecture to model synthetic pathways in a generative manner. MoleculeChef is one of the earliest neural models of this kind. Instead of synthetic paths, it tries to encode and decode a bag of starting reactants from purchasable building blocks. Following an enumeration of the one-step synthetic pathways, the “best” molecule from the products is chosen as the output. DoGs represents synthetic pathways as directed acyclic graphs and serializes the graph construction process as a sequence of actions that includes node addition, reactant selection, and node connection. DoGs use recurrent neural networks to generate action sequences from latent codes, enabling de novo molecular generation by drawing samples from the latent space. SynNet (Gao et al., 2021) can also be seen as this kind, in which the molecular fingerprinting algorithm is the encoder, and neural networks are used as a decoder to generate synthetic paths.
@@
Limitations: While the synthesizable molecular design methods have advanced significantly, they are still in the proof-of-concept stage and far from mature. The efficiency of combinatorial optimization often falls short, as these algorithms edit molecules locally, thus being inefficient to explore the vast chemical space. Reinforcement learning techniques depend heavily on querying oracle functions, which can hinder their practical application. MoleculeChef (Bradshaw et al., 2019) primarily explores one-step pathways due to the computational impracticality of enumerating an exponentially growing number of candidates. Models like PGFS (Gottipati et al., 2020) and REACTOR (Horwood & Noutahi, 2020) are limited to generating only linear synthetic pathways. In addition, the capacity of these models to generalize and construct complex synthetic paths is yet to be convincingly demonstrated. Models such as DoGs (Bradshaw et al., 2020) and SynNet (Gao et al., 2021), theoretically capable of encompassing a complete chemical space, have not been empirically validated for their ability to adapt to novel, unseen chemicals. Additionally, these models are generally less adept at generating convergent pathways or more intricate molecular structures, indicating a need for further refinement in their design and functionality.
@@
Figure 2: Top: Simulation of the synthetic plan $R_{2}(R_{1}(B_{1},B_{2}),B_{3})$ with a memory stack to store intermediate molecules. In the first two steps, building blocks $B_{1}$ and $B_{2}$ are pushed onto the stack. In the third step, $B_{1}$ and $B_{2}$ are popped from the stack and used as input for reaction $R_{1}$, leading the intermediate reaction product $I$, which is then added back onto the stack. Subsequently, in the fourth step, building block $B_{3}$ is pushed onto the stack. Finally, reaction $R_{2}$ is applied to $I$ and $B_{3}$, leading to the final product $Y$. Bottom: The postfix notation of the synthetic pathway is $[B_{1},B_{2},R_{1},B_{3},R_{2}]$. Each symbol in the notation corresponds to one step in the synthesis.
@@ 3.1 Preliminaries
A chemical space ${\mathcal{C}}$ is a set of molecules generated by a finite set of building block molecules =${\mathcal{B}}=\{B_{1},B_{2},\ldots B_{N}\}\subseteq{\mathcal{C}}$ and a set of reaction rules =${\mathcal{R}}=\{R_{1},R_{2},\ldots R_{M}\}$. Formally, a reaction rule maps reactants to a reaction product. For example, a reaction rule $R$ with two reactants can be denoted by:
@@
Following the above example, the instruction sequence for $R_{2}(R_{1}(B_{1},B_{2}),B_{3})$ is: (1) push $B_{1}$; (2) push $B_{2}$; (3) apply $R_{1}$; (4) push $B_{3}$; (5) apply $R_{2}$, which can be simplified as the postfix notation: $[B_{1},B_{2},R_{1},B_{3},R_{2}]$ (Figure 2).
The synthetic process can be parsed to an abstract syntax tree (AST), which is equivalent to the synthetic tree used in previous work and the postfix notation is a post-order traversal sequence of the AST. Although it has connections with previous methods, the proposed postfix notation has multiple advantages. First, the postfix notation is capable of representing branched synthesis pathways, where reactions can be applied to more than one intermediate deposited in the stack. In contrast, previous methods (Gottipati et al., 2020; Horwood & Noutahi, 2020) are limited to linear synthetic pathways, where intermediates can only react with building blocks. In addition, the postfix notation is a simple representation comprising only building blocks and reactions, whereas previous synthetic tree-based representations require extra symbols, such as node addition and edge addition, to construct a synthetic plan. Lastly, the postfix notation allows reactions with multiple (more than two) molecule inputs, while the previous representation (Gao et al., 2021) is incompatible with such complex reaction rules. Therefore, the postfix notation can be immediately applied to chemical spaces that include reactions with reagent specification or with multiple reactants, showing the potential for more complex settings in future work.
@@
Figure 3: (a) Overall architecture of the network. A graph transformer encodes the input molecular graph into atom embeddings. A transformer decoder autoregressively decodes the postfix notation, with atom embeddings as the context. (b) Architecture of the prediction network. First, a classifier predicts the token type from the token embedding. If the predicted token type is [BB] (building block), the token embedding is used to predict the building block fingerprint to retrieve building blocks. If the token type is [RXN] (reaction), the token embedding is used to predict the type of reaction. Otherwise, [END] marks the end of the postfix notation.
@@ 3.2 Model
A molecule is represented by a graph with a set of atoms and bonds. The graph encoder network first converts each atom into initial embeddings according to the atomic number. Then a stack of graph transformer layers (Ying et al., 2021), which take edges into account by adding a bond type-dependent learnable bias term to the query-key product matrix, performs multi-head self-attention on the atoms. Finally, the output per-atom embeddings are used for decoding.
@@
A postfix notation of synthesis is a sequence that contains four types of tokens: building block tokens, reaction tokens, a start token [START], and an end token [END]. A building block token is attached with the fingerprint of the building block molecule, denoted by [BB,${\bm{p}}$], where ${\bm{p}}\in\{0,1\}^{256}$ is the Morgan fingerprint of length 256 and radius 2 (Morgan, 1965). Building block tokens are converted into embeddings by an MLP. A reaction token is the index $r$ of the reaction denoted by [RXN,$r$]. The reaction tokens and the start token are converted into embeddings by looking up from the embedding table. Sinusoidal positional encodings (Vaswani et al., 2017) that indicate the order of tokens are added to the embeddings. In summary, the initial embedding of the $i$-th token $T_{i}$ is defined as:
@@
Next, a stack of standard transformer layers (Vaswani et al., 2017) takes the sequence embeddings along with the molecular graph embeddings as input, and then predicts the next token using the embeddings from the last layer ${\bm{h}}_{i}$ for each position of the sequence. The prediction network first predicts the type of the next token:
@@ 3.3 Training
We use the standard technique to train transformer decoders in parallel (Vaswani et al., 2017) — feeding the complete sequence into the decoder, offseting the output embeddings by one position, and applying causal masking.
@@
Dataset Method SuccessRecons.Sim.(Morgan) Sim.(Scaffold) Sim.(Gobbi) Test Set SynNet 84.1% 10.7% 0.4575 0.5109 0.3465 Proposed 97.5% 28.4% 0.7167 0.7791 0.7273 ChEMBL SynNet 85.0% 5.4% 0.4270 0.4174 0.2678 Proposed 98.8% 13.3% 0.5978 0.5869 0.5570

### v2

@@ Abstract
Discovering new drug molecules is a pivotal yet challenging process due to the near-infinitely large chemical space and notorious demands on time and resources. Numerous generative models have recently been introduced to accelerate the drug discovery process, but their progression to experimental validation remains limited, largely due to a lack of consideration for synthetic accessibility in practical settings. In this work, we introduce ChemProjector, a novel framework that is capable of generating new chemical structures while ensuring synthetic accessibility. Specifically, we introduce a postfix notation of synthetic pathways to represent molecules in chemical space. Then, we design a transformer-based model to translate molecular graphs into postfix notations of synthesis. We highlight the model’s ability to: (a) perform bottom-up synthesis planning more accurately, (b) generate structurally similar, synthesizable analogs for unsynthesizable molecules proposed by generative models with their properties preserved, and (c) explore the local synthesizable chemical space around hit molecules. Code is available at https://github.com/luost26/ChemProjector.
@@ 1 Introduction
Designing novel drug molecules, a process that aims to identify molecular structures with desired properties, is known to be notoriously complex, time- and resource-consuming (wouters2020estimated). Methods that could accelerate the discovery process are thus of significant interest in medicinal science and the pharmaceutical industry. Historically, two principal computational methodologies have been employed in drug design: virtual screening (shoichet2004virtual) and de novo design (hartenfeller2011novo). Virtual screening involves predicting properties for a predefined set of molecules, known as virtual libraries, and identifying the most promising candidates for subsequent validation. These libraries are usually commercial catalogs (make-on-demand compounds) from chemical vendors (lyu2019ultra) and generally consist of molecules with assured synthesizability (enamine). However, screening large numbers of molecules is a time-intensive process. Moreover, even libraries on the scale of billions present only a tiny fraction of the theoretical chemical space (Figure 1a), which is nearly infinite in the number of potential synthesizable small molecules (bohacek1996art). This means that many potentially valuable molecules might not be included in these libraries. Additionally, as the size of the library increases, so does the cost of screening, which in turn limits the effectiveness and efficiency of virtual screening in the drug discovery process.
@@
De novo molecular design methods, unlike virtual screening, dynamically create molecules tailored to specific objectives like protein binding and bioactivity. Early methods of this kind typically adopt combinatorial optimization (venkatasubramanian1995evolutionary; jensen2019graph; brown2019guacamol), while recently, neural network enhanced optimization (you2018graph; fu2021differentiable) and generative modeling (gomez2018automatic; jin2018junction) have become popular. These methods aren’t constrained by the limitations of virtual libraries, enabling the efficient exploration of novel chemical entities (see Figure 1c). However, a notable defect with most existing generative models is their tendency to propose synthetically infeasible molecules (Figure 1c) (gao2020synthesizability), a concern particularly pronounced in structure-based drug design, an area of increasing interest (bohacek1996art; luo20213sbdd; fu2022reinforced). This issue stems from the inherent ignorance of synthetic accessibility in formulations of most existing generative algorithms, in which molecules are constructed using SMILES strings or atoms and motifs as fundamental units (gao2022sample). This often leads to chemical structures that fall outside synthesizable chemical spaces, posing practical challenges in the experimental validation as well as production of such designs.
@@
Incorporating synthesizability as an additional optimization criterion is a common strategy, yet it introduces complex challenges. Synthesizability is difficult to define precisely due to its sensitivity to subtle structural differences, reaction selectivity, and building block availability inherent in chemical reactions (gao2020synthesizability). While notable efforts have been made to estimate synthetic accessibility (ertl2009estimation; thakkar2021retrosynthetic; vorvsilak2020syba), reliable quantification of synthetic accessibility is still a far goal. A more effective strategy involves constraining the design process to only consider molecules that are synthesizable, ensuring that algorithms focus exclusively on viable structures. Based on this idea, several models have been introduced (bradshaw2019moleculechef; bradshaw2020barking; gao2021synnet; swanson2023generative), though these models have notable limitations, such as poor coverage of synthesizable chemical space or inefficiency in optimizing the properties, as further discussed in Section 2.
@@
To address the recognized challenges, we introduce a framework that learns to “project” molecules into synthesizable chemical spaces (Figure 1d). Given molecules without synthesizability guarantee, our model is able to “fix” them by identifying synthesizable analogs that are structurally similar, thus preserving the structural factors that contributed to the desired properties (janela2022simple). We represent molecules by a scalable linear representation for synthetic pathways based on postfix notations. By generating synthetic pathways in the form of postfix notations rather than molecular graphs or strings, we can ensure the designed molecules are derivable from purchasable chemical building blocks and expert-defined chemical reaction rules, thereby guaranteeing the synthesizability of resulting molecules. We adopt the transformer architecture (vaswani2017attention), with an encoder for molecular graphs and a decoder to generate postfix notations of synthesis.
@@ 2 Related Work
Combinatorial optimization algorithms use a set of chemical building blocks and reaction templates to make the chemical space combinatorial, and then apply combinatorial optimization algorithms, such as genetic algorithms (gao2021synnet), Monte Carlo tree search (swanson2023generative), and Bayesian optimization (korovina2020chembo), to search for desired molecules (Figure 1b). Early representative work before the advent of molecular deep learning includes SYNOPSIS (vinkers2003synopsis) and DOGS (hartenfeller2012dogs), which perform virtual reactions to generate candidate molecules and select candidates according to scoring functions. In recent years, advances in deep learning for chemical reactions (coley2019reactivity; schwaller2019molecular) have inspired methods such as MoleculeChef (bradshaw2019moleculechef), ChemBO (korovina2020chembo), and DoGs (bradshaw2020barking) that use neural networks to predict reaction outcomes in a “template-free” manner.
@@
Reinforcement learning approaches, exemplified by PGFS (gottipati2020pgfs) and REACTOR (horwood2020reactor), formulate the forward synthesis process as Markov decision processes and train agents to explore the chemical space under the guidance of reward functions that quantify desired properties. The agent is provided with a set of reactants and reaction templates and learns to synthesize desirable molecules with them. As the agent is aware of the objective, its search process is directed and expected to be more efficient.
@@
Autoencoder-based methods, represented by MoleculeChef (bradshaw2019moleculechef) and DoGs (bradshaw2020barking), leverage the autoencoder architecture to model synthetic pathways in a generative manner. MoleculeChef is one of the earliest neural models of this kind. Instead of synthetic paths, it tries to encode and decode a bag of starting reactants from purchasable building blocks. Following an enumeration of the one-step synthetic pathways, the “best” molecule from the products is chosen as the output. DoGs represents synthetic pathways as directed acyclic graphs and serializes the graph construction process as a sequence of actions that includes node addition, reactant selection, and node connection. DoGs use recurrent neural networks to generate action sequences from latent codes, enabling de novo molecular generation by drawing samples from the latent space. SynNet (gao2021synnet) can also be seen as this kind, in which the molecular fingerprinting algorithm is the encoder, and neural networks are used as a decoder to generate synthetic paths.
@@
Limitations: While the synthesizable molecular design methods have advanced significantly, they are still in the proof-of-concept stage and far from mature. The efficiency of combinatorial optimization often falls short, as these algorithms edit molecules locally, thus being inefficient to explore the vast chemical space. Reinforcement learning techniques depend heavily on querying oracle functions, which can hinder their practical application. MoleculeChef (bradshaw2019moleculechef) primarily explores one-step pathways due to the computational impracticality of enumerating an exponentially growing number of candidates. Models like PGFS (gottipati2020pgfs) and REACTOR (horwood2020reactor) are limited to generating only linear synthetic pathways. In addition, the capacity of these models to generalize and construct complex synthetic paths is yet to be convincingly demonstrated. Models such as DoGs (bradshaw2020barking) and SynNet (gao2021synnet), theoretically capable of encompassing a complete chemical space, have not been empirically validated for their ability to adapt to novel, unseen chemicals. Additionally, these models are generally less adept at generating convergent pathways or more intricate molecular structures, indicating a need for further refinement in their design and functionality.
@@
Figure 2: Top: Simulation of the synthetic plan $R_{2}(R_{1}(B_{1},B_{2}),B_{3})$ with a memory stack to store intermediate molecules. In the first two steps, building blocks $B_{1}$ and $B_{2}$ are pushed onto the stack. In the third step, $B_{1}$ and $B_{2}$ are popped from the stack and used as input for reaction $R_{1}$, leading the intermediate reaction product II, which is then added back onto the stack. Subsequently, in the fourth step, building block $B_{3}$ is pushed onto the stack. Finally, reaction $R_{2}$ is applied to II and $B_{3}$, leading to the final product $Y$. Bottom: The postfix notation of the synthetic pathway is $B_{1},B_{2},R_{1},B_{3},R_{2}]$. Each symbol in the notation corresponds to one step in the synthesis.
@@ 3.1 Preliminaries
A chemical space $\mathcal{C}$} is a set of molecules generated by a finite set of building block molecules ${\mathcal{B}}=\{B_{1},B_{2},\ldots B_{N}\}\subseteq{\mathcal{C}}$ and a set of reaction rules ${\mathcal{R}}=\{R_{1},R_{2},\ldots R_{M}\}$. Formally, a reaction rule maps reactants to a reaction product. For example, a reaction rule $R$ with two reactants can be denoted by:
@@
Following the above example, the instruction sequence for $R_{2}(R_{1}(B_{1},B_{2}),B_{3})$ is: (1) push $B_{1}$; (2) push $B_{2}$; (3) apply $R_{1}$; (4) push $B_{3}$; (5) apply $R_{2}$, which can be simplified as the postfix notation: $B_{1},B_{2},R_{1},B_{3},R_{2}]$ (Figure 2).
The synthetic process can be parsed to an abstract syntax tree (AST), which is equivalent to the synthetic tree used in previous work and the postfix notation is a post-order traversal sequence of the AST. Although it has connections with previous methods, the proposed postfix notation has multiple advantages. First, the postfix notation is capable of representing branched synthesis pathways, where reactions can be applied to more than one intermediate deposited in the stack. In contrast, previous methods (gottipati2020pgfs; horwood2020reactor) are limited to linear synthetic pathways, where intermediates can only react with building blocks. In addition, the postfix notation is a simple representation comprising only building blocks and reactions, whereas previous synthetic tree-based representations require extra symbols, such as node addition and edge addition, to construct a synthetic plan. Lastly, the postfix notation allows reactions with multiple (more than two) molecule inputs, while the previous representation (gao2021synnet) is incompatible with such complex reaction rules. Therefore, the postfix notation can be immediately applied to chemical spaces that include reactions with reagent specification or with multiple reactants, showing the potential for more complex settings in future work.
@@
Figure 3: (a) Overall architecture of the network. A graph transformer encodes the input molecular graph into atom embeddings. A transformer decoder autoregressively decodes the postfix notation, with atom embeddings as the context. (b) Architecture of the prediction network. First, a classifier predicts the token type from the token embedding. If the predicted token type is [$B$] (building block), the token embedding is used to predict the building block fingerprint to retrieve building blocks. If the token type is [RXN] (reaction), the token embedding is used to predict the type of reaction. Otherwise, [END] marks the end of the postfix notation.
@@ 3.2 Model
A molecule is represented by a graph with a set of atoms and bonds. The graph encoder network first converts each atom into initial embeddings according to the atomic number. Then a stack of graph transformer layers (ying2021graphormer), which take edges into account by adding a bond type-dependent learnable bias term to the query-key product matrix, performs multi-head self-attention on the atoms. Finally, the output per-atom embeddings are used for decoding.
@@
[…]
@@
Next, a stack of standard transformer layers (vaswani2017attention) takes the sequence embeddings along with the molecular graph embeddings as input, and then predicts the next token using the embeddings from the last layer ${\bm{h}}_{i}$ for each position of the sequence. The prediction network first predicts the type of the next token:
@@ 3.3 Training
We use the standard technique to train transformer decoders in parallel (vaswani2017attention) — feeding the complete sequence into the decoder, offseting the output embeddings by one position, and applying causal masking.
@@
Dataset Method Success% Recons.% Sim.(Morgan) Sim.(Scaffold) Sim.(Gobbi) Test Set SynNet 84.1% 10.7% 0.4575 0.5109 0.3465 ChemProjector 97.5% 28.4% 0.7167 0.7791 0.7273 ChEMBL SynNet 85.0% 5.4% 0.4270 0.4174 0.2678 ChemProjector 98.8% 13.3% 0.5978 0.5869 0.5570